OpenStudy (anonymous):

Choose the polynomial that is written in standard form. A. −3x5y2 + 4x3y + 10x2 B. −8xy2 + 4x4y2 + 3x3 C. x4y2 + 4x3y5 + 10x4 D. x6y2 + 4x3y8 + 10x7 Now this one looks a little confusing

OpenStudy (anonymous):

first one starting to read from left to right the powers must reduce

OpenStudy (anonymous):

So the first one is written in standard form? what about C though

OpenStudy (anonymous):

in C the second term: the power of y rises from 2 to 5 which is not standard

OpenStudy (anonymous):

Can I have you to help me on the rest Looks like you are really smart

OpenStudy (anonymous):

u can count on me

OpenStudy (anonymous):

Choose the correct sum of the polynomials (4x3 − 2x − 9) + (2x3 + 5x + 3). a, 6x3 − 3x − 6 b. 2x3 − 7x − 12 C. 6x3 + 3x − 6 D. 2x3 + 3x − 3

OpenStudy (anonymous):

I got B for this one

OpenStudy (anonymous):

why? just wanna know where u make mistake

OpenStudy (anonymous):

It says find the sum Dont i just add the numbers together or? like 2x + 5x = 7x

OpenStudy (anonymous):

yes but u must add the coefficients of the same power of x right? then it would be c

OpenStudy (anonymous):

Oh okay i get what u mean

OpenStudy (anonymous):

Choose the correct simplification of the expression (4x3y2z4)(2x3y4z3). A.8x9y8z12 B. 6x6y6z7 C. 8x6y6z7 D. 6x9y8z12

OpenStudy (anonymous):

Ik its either A or C

OpenStudy (anonymous):

Its A cause u multiply right

OpenStudy (anonymous):

u confused me which are the powers?

OpenStudy (anonymous):

could u use the equation editor?

OpenStudy (anonymous):

4x times 2x= 8x then 3y times 3y is 9y then 2z times 4z is 8z then 4 times 3 is 12 And those answers are all in A

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

well none of them were powers!

OpenStudy (anonymous):

and BTW uor right its A

OpenStudy (anonymous):

and BTW ur right its A

OpenStudy (anonymous):

Choose the correct simplification of the expression \[(5a ^{7}b)^{2} \]

OpenStudy (anonymous):

what do u think?

OpenStudy (anonymous):

\[A. 25^{14}b^{2} \]

OpenStudy (anonymous):

correct but I think u missed sth.

OpenStudy (anonymous):

I did?

OpenStudy (anonymous):

Oh i forgot the a im sorry

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Choose the correct simplification of the expression (5xy5)2(y3)4.\[(5xy^{5})^{2}(y^3)4.\]

OpenStudy (anonymous):

Im sorry the for is a exponent

OpenStudy (anonymous):

four*

OpenStudy (anonymous):

ok \[25x ^{2}y ^{18}\]

OpenStudy (anonymous):

Choose the correct simplification of the expression \[ab^{9}b^{10}\a^{2}b^{7}\]

OpenStudy (anonymous):

Its a fraction

OpenStudy (anonymous):

whats that backslash?

OpenStudy (anonymous):

Its fraction bar

OpenStudy (anonymous):

\[\frac{ b ^{12} }{ a }\]

OpenStudy (anonymous):

a11b17 1 over a to the 11th power times b to the 17th power 1 over a to the 7th power times b to the 3rd power a7b3

OpenStudy (anonymous):

these are my choises

OpenStudy (anonymous):

is fraction bar is a dividing operator? u see I don't know all of the math expressions in english!

OpenStudy (anonymous):

\[a^{11}b^{17}\]

OpenStudy (anonymous):

\[\frac{ 1 }{ a^{11}b^{17} }\]

OpenStudy (anonymous):

\[\frac{ 1 }{ a^{7}b^{3} }\]

OpenStudy (anonymous):

\[a^{7}b^{3}\]

OpenStudy (anonymous):

ok Thats better

OpenStudy (anonymous):

lol now those are the choices

OpenStudy (anonymous):

none of them are correct

OpenStudy (anonymous):

9+2 is 11 10+7 is 17 its either A or B correct?

OpenStudy (anonymous):

@ankit042

OpenStudy (anonymous):

\[\frac{ a^9b^{10} }{ a^{2}b^{7} }\]

OpenStudy (anonymous):

Thats what it looks like

OpenStudy (ankit042):

this is simple a^x/a^y = a^(x-y) apply this rule for a and b you will get your answer

OpenStudy (anonymous):

I got B

OpenStudy (ankit042):

which is?

OpenStudy (anonymous):

\[\frac{ 1 }{ a^{11}b^{17} }\]